Jacob and Kayla are selling cheesecake for a school fundraiser. Customers can buy pecan cheesecake. Jacob sold 2 pecan cheesecake and 4 apple cheesecake for a total of 96$. Kayla Sold 8 pecan cheesecake and 4 apple cheesecake for a total of 156$ . What is the cost each of one pecan cheesecake and one apple cheesecake?
Answer by greenestamps(13200)



The difference between the two sales is 6 additional pecan cheesecakes, for a cost difference of $60. So each pecan cheesecake costs $60/6 = $10.

Then using either of the sales and the $10 cost of each pecan cheesecake to determine the cost of each apple cheesecake.

Algebraically....

(1) 2p+4a = 96
(2) 8p+4a = 156

Subtract (1) from (2):
6p = 60
p=60/6=10

Substitute p=10 in either (1) or (2):
20+4a = 96
4a = 76
a = 76/4 = 19

ANSWERS: $10 for each pecan cheesecake; $19 for each apple cheesecake